Circular No. 5748

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 H. U. Zimmermann, Max-Planck-Institut für Extraterrestrische Physik

(MPE); W. Lewin and E. Magnier, Massachusetts Institute of Technology; P. Predehl, G. Hasinger, W. Pietsch, B. Aschenbach and J. Trümper, MPE; G. Fabbiano, Center for Astrophysics; J. van Paradijs, University of Amsterdam; L. Lubin, Princeton University; and R. Petre, Goddard Space Flight Center, communicate: “We observed SN 1993J with ROSAT on Apr. 3.4 UT. In a 2826-s exposure using the PSPC detector the supernova shows up as a bright new source within 10′′ of the optical position as given on IAUC 5731. At this time only raw telemetry data are available, so a calibrated x-ray position has to await the final attitude solution. The xray supernova is located ∼ 1′ northwest of a bright x-ray source that was already detected by Einstein and ROSAT (source X-6 in the paper by Fabbiano 1988, Ap.J. 325, 544). A first estimate of the intensity shows that the supernova is ∼ 1.7 times stronger than X-6 in soft x-rays.”

A. C. Porter and L. A. Wells, Kitt Peak National Observatory, report: “A high-resolution (0.08 nm) spectrum taken by V. C. Rubin at the KPNO 4-m telescope on Mar. 31.19 UT shows a low-contrast (EW = 0.014 nm) feature centered at 656.5 nm. Interpreted as Hα, this corresponds to a radial velocity of ∼ 150 km/s relative to the systemic value for M81. Several weak (EW = 0.003 nm) absorption features are visible in the region from 645 to 670 nm. The 656.5-nm line is almost certainly the same one reported by Filippenko and Matheson on IAUC 5740. At high resolution, it shows a sharp blue edge, a sloping red side and a probable weak absorption feature at its center. Such a line profile is consistent with emission from a dusty or thick shell or sphere, perhaps from a wind phase late in the progenitor’s lifetime. The line’s FWHM is 1.2 nm, corresponding to a shell velocity of 300 km/s. Such a shell should be consumed by the supernova shock within a few weeks or months, causing the disappearance of this line over the same timescale, as in the more dramatic example of SN 1984E (cf. IAUC 5742). High-resolution observations of this and the other faint lines reported on IAUC 5740 are strongly urged to follow the evolution of this feature.”

A. V. Filippenko, University of California at Berkeley, reports that M. Lehnert (Lawrence Livermore National Laboratory) and P. Butler (also of Berkeley) obtained a high-dispersion spectrum of SN 1993J on Mar. 31 UT with the Shane 3-m reflector at the Lick Observatory. The resolution of ∼ 0.015 nm reveals four strong components and several weak ones in the Na I D interstellar absorption lines.

1993 April 5 Brian G. Marsden